# TOAD: Task-Oriented Automatic Dialogs with Diverse Response Styles

### v2

@@ Abstract
In light of recent advances in large language models (LLMs), the expectations for the next generation of virtual assistants include enhanced naturalness and adaptability across diverse usage scenarios. However, the creation of high-quality annotated data for Task-Oriented Dialog (TOD) is recognized to be slow and costly. To address these challenges, we introduce Task-Oriented Automatic Dialogs (TOAD), a novel and scalable TOD dataset along with its automatic generation pipeline. The TOAD dataset simulates realistic app context interaction and provide a variety of system response style options. Two aspects of system response styles are considered, verbosity level and users’ expression mirroring. We benchmark TOAD on two response generation tasks and the results show that modeling more verbose or responses without user expression mirroring is more challenging.  [^1] The data and code are avaliable at https://github.com/williamLyh/toad.
@@ 2.1 Task-Oriented Dialog Dataset
To our best knowledge, TOAD stands out as the first highly automatically generated TOD dataset. All previous datasets relied on human annotations or paraphrasing, which makes the collection process costly and time-consuming. Moreover, our dataset also admits all of the unique properties listed in Section 1.
@@ 2.2 Dialog Distillation
Recent advancements in LLMs have demonstrated their improved ability to follow human instructions and generate fluent conversation utterances for a range of understanding and generation tasks. A list of recent works embrace this concept, extracting knowledge and dialog from LLMs to train student or in-domain models. SODA (Kim et al., 2023) distilled social interactions dialogue by contextualizing commonsense knowledge from a social knowledge graph. PLACES (Chen et al., 2023) generated a multi-party social dialog by prompting the OPT (Zhang et al., 2022) model with topic, facts, and dialog as in-context learning examples. MathDial (Macina et al., 2023) distilled teacher-student math reasoning tutoring dialog through pairing human teachers with a LLM. There are also works that distil multi-turn open-domain instructional conversations, such as Baize (Xu et al., 2023) and UltraChat (Ding et al., 2023). Our TOAD represents the first work in distilling TOD dialog with multiple response style options, which requires more nuanced control over the conversation content.
@@ 3.1 Persona-Grounded Context Generation
Data diversity is important to the model’s robustness and generalization. However, LLM-generated data given the same input prompt often lacks diversity. To address the two issues mentioned above, we develop a persona-grounded context generation pipeline, which combines sampling from external data sources and chain-of-thought generation (Wei et al., 2022).
@@
  Compound restaurant_booking_reserve_table(restaurant ="French Brasserie", time="8:00 PM"),
I would like to reserve a table at the French Brasserie for 8:00 PM. Also, can you help me find a hotel in Las Vegas?
@@
  Compositional weather_get_weather(date=get_calendar_events( name="Art Class").calendar_events_check(date).date)
What’s the weather like on the day of my Art Class event?
  Self-correction get_movie_time(movie_name="Fast & Furious Presents: Hobbs & Shaw", location="Miami").self_correction( location="Houston")
Could you find the showtimes for Fast & Furious Presents: Hobbs & Shaw in Miami? Actually, make that Houston instead.
  Complex referral get_alarms(ordered_by="time", index=0).check(time)
What’s the time for my earliest alarm?
@@ 3.4 Dataset Quality Control
Following the generation pipeline, we filter the dataset to ensure its quality, which involved two filtering conditions: 1) Removing datapoints with utterances with unnatural formats, such as canonical datetime expressions like ‘2025-12-16’. 2) Removing datapoints that have missing slot values in the corresponding utterances. This filtering process resulted in the removal of 2.2% of the data, yielding a final dataset size of 8087 dialogs.
@@ 9 Ethics and Risks
Term of Use
The dataset and generation pipeline of TOAD is released under CC-BY 4.0 license.
Persona Bias in TOAD
The persona, e.g., ethnicity, gender, and occupation distributions of the TOAD dataset are elaborated in detail in Appendix E.
@@ Appendix E Persona and App Context Statistics
In TOAD, we simulate 300 personas with diverse backgrounds, with one persona randomly assigned to each dialog as the user’s information source. The gender distribution of the simulated personas is approximately 52% male and 48% female.
@@ Appendix G Plot and Dialog Examples
{’today’:’Wed 2025-04-16’,
’alarms’:[{’time’:’07:00’, ’name’:’Morning Workout’, ’if_repeat’:True}, {’time’:’18:30’, ’name’:’Family Dinner’, ’if_repeat’:True}]}
User@: [get_alarms(ordered_by="time", index=0).alarms_delete(name), get_alarms(ordered_by="time", index=1).alarms_check(name)]
System@: [alarms_ask_confirm(get_alarms(name="Morning Workout").delete(name)), alarms_inform_result(name="Family Dinner")]
User@: alarms_confirm()
System@: alarms_notify_done(get_alarms(name="Morning Workout").delete(name))
User@: Can you delete my earliest alarm and check the next one?
System@: {
HV M: Are you sure you want to delete the Morning Workout alarm? Your next alarm is Family Dinner at 18:30.
HV no_M: Do you confirm the deletion of your earliest alarm, Morning Workout’? The following alarm is Family Dinner at 18:30.
MV M: You want to delete the Morning Workout alarm, correct? Your next one is Family Dinner at 18:30.
MV no_M: Confirm to delete this alarm? The next one is at 18:30 for Family Dinner’.
LV M: Delete Morning Workout’? Next: Family Dinner 18:30.
LV no_M: Confirm deletion? Next: 18:30 Family Dinner’. }
User@: Yes, go ahead and delete it.
System@: {
HV M: Your Morning Workout alarm has been deleted.
HV no_M: The earliest alarm has been deleted.
MV M: Morning Workout alarm is deleted.
MV no_M: Its deleted.
LV M: Deleted.
LV no_M: Done.}
{’today’:’Mon 2026-12-28’}
User@: purchase_tickets(movie_name="Joker", date="Tue 2026-12-29", cinema_name="Regal Cinemas", ticket_quantity=1, movie_format="3d")
System@: request_infomation(showtime)
User@: inform_infomation(showtime="20:00")
System@: ask_confirm(purchase_tickets(movie_name="Joker", ticket_quantity=1, date="Tue 2026-12-29", showtime="20:00", movie_format="3d", cinema_name="Regal Cinemas"))
User@: confirm()
System@: notify_done(purchase_tickets(movie_name="Joker", ticket_quantity=1, date="Tue 2026-12-29", showtime="20:00", movie_format="3d", cinema_name="Regal Cinemas"))
User@: I’d like to buy a 3D ticket for Joker at Regal Cinemas tomorrow, please.
System@: {
HV M: What time would you like to watch Joker at Regal Cinemas tomorrow?
HV no_M: What showtime do you prefer?
MV M: When would you like to see it?
MV no_M: When is the preferred time?
LV M: Showtime?
LV no_M: Time? }
User@: At 20:00.
System@: {
HV M: So, youd like to purchase a 3D ticket for Joker at Regal Cinemas on Tue 2026-12-29 at 20:00. Is that correct?
HV no_M: Just to confirm, you want to buy a 3D ticket for the movie Joker at Regal Cinemas on Tue 2026-12-29 at 20:00, right?
MV M: Youre buying a 3D ticket for Joker at Regal Cinemas tomorrow at 20:00. Correct?
MV no_M: So, youre purchasing a ticket for Joker in 3D at Regal Cinemas tomorrow at 20:00. Is that right?
LV M: Confirm purchase?
LV no_M: Confirm? }
User@: Yes, that’s correct.
System@: {
HV M: Your 3D ticket for Joker at Regal Cinemas on Tue 2026-12-29 at 20:00 has been purchased.
HV no_M: The ticket has been successfully purchased.
MV M: Ive bought your ticket for that movie at that time.
MV no_M: Its done, the ticket is bought.
LV M: Done.
LV no_M: Purchased. }
@@ Appendix H Dialog Generator template
Prompt template for user utterance generation
You are a smartphone user and you are testing your virtual assistant on your phone by \
engaging in a multi-turn conversations with it.
Here is your personal introduction: {{ user_intro }}
Instructions:
1. You need to communicate with the assistant following the guidance of "actions".
2. Based on the introduction, think about what speech habit you should have and \
communicate with this pattern.
3. The "message" should have minimum words possible.
4. You must return in JSON format, following the provided examples.
The context information is: {{ context }}.
You are using these apps: {{ situation }}.
Begin conversation (you are identified as "user").
user: {"actions": ["hello()"], "message": "Hi."}
assistant: {"actions": ["offer_help()"], "message": "Hello, how can I help?"}
{if dialog_history | length>0 %}{{ dialog_history }}{endif %}
The "actions" you need to follow is {{ action[cur_turn] }}. What are you going to say next? \
{{ user_style_instruction }}
user:\
@@ 2. Based on the introduction, think about what speech habit you should have and \
Instructions:
1. You are a virtual assistant. Your goal is to assist the user to accomplish their goal.
2. Your responses should strictly follow the given actions and be helpful, natural, \
professional and concise.
3. Your response should strictly follow the corresponding "actions".
The user is interacting with these apps: {{ situation }}.
Style instruction:
verbosity_low’: Your response must only have a couple of words, such as "when", \
"how long" or "done".
verbosity_mid’: Your response should be a concise but complete sentence and must \
replace the nouns or noun phrases mentioned by user with pronouns, such as "it", "that" and "its".
verbosity_high’: Your response should use full expressions with all the details.
mirroring’: Your response should use the users noun phrase or verb expressions when possible.
no_mirroring’: Ignore all previous dialog. Do not affect by user expression.
summary’: Your response should be a brief report of the given summary.
Response Example:
assistant actions: {"verbosity_low mirroring": ["notify_done()"], \
"verbosity_low no_mirroring": ["notify_done()"], \
"verbosity_mid mirroring": ["notify_done(operation_on_device(\
operation="turn_off", device="heating", home_space="bedroom"))"], \
"verbosity_mid no_mirroring": ["notify_done(operation_on_device(\
operation="turn_off", device="heating", home_space="bedroom"))"], \
"verbosity_high mirroring": ["notify_done(operation_on_device(\
operation="turn_off", device="heating", home_space="bedroom"))"], \
"verbosity_high no_mirroring": ["notify_done(operation_on_device(\
operation="turn_off", device="heating", home_space="bedroom"))"]}
assistant: {"verbosity_low mirroring": "Turned off.", \
"verbosity_low no_mirroring": "Done.", \
"verbosity_mid mirroring": "I have turned off the heating in that room.", \
"verbosity_mid no_mirroring": "I have turned it off in that room.", \
"verbosity_high mirroring": "I have turned off the bedroom heating.", \
"verbosity_high no_mirroring": "Sure, I have turned off the heating in the bedroom."}
Conversation history:
{if dialog_history | length>0 %}{{ dialog_history }}{endif %}
New turn:
You should return in JSON format with 6 keys: ["verbosity_low mirroring", \
"verbosity_low no_mirroring", "verbosity_mid mirroring", "verbosity_mid no_mirroring", \
"verbosity_high mirroring", "verbosity_high no_mirroring"].
assistant actions: {{ action[cur_turn] }}.
assistant: \

### v3

@@ Abstract
In light of recent advances in large language models (LLMs), the expectations for the next generation of virtual assistants include enhanced naturalness and adaptability across diverse usage scenarios. However, the creation of high-quality annotated data for Task-Oriented Dialog (TOD) is recognized to be slow and costly. To address these challenges, we introduce Task-Oriented Automatic Dialogs (TOAD), a novel and scalable TOD dataset along with its automatic generation pipeline. The TOAD dataset simulates realistic app context interaction and provide a variety of system response style options. Two aspects of system response styles are considered, verbosity level and users’ expression mirroring. We benchmark TOAD on two response generation tasks, and the results show that modeling more verbose responses or responses without user expression mirroring is more challenging. [^1] The data and code are available at https://github.com/apple/ml-toad.
@@ 2.1 Task-Oriented Dialog Dataset
To our best knowledge, TOAD stands out as the first highly automatically generated TOD dataset. A concurrent work, LUCID (Stacey et al., 2024), also adopts a similar LLM-driven approach for dialog generation. All previous datasets relied on human annotations or paraphrasing, which makes the collection process costly and time-consuming. Moreover, our dataset also admits all of the unique properties listed in Section 1.
@@ 2.2 Dialog Distillation
Recent advancements in LLMs have demonstrated their improved ability to follow human instructions and generate fluent conversation utterances for a range of understanding and generation tasks. A list of recent works embrace this concept, extracting knowledge and dialog from LLMs to train student or in-domain models. SODA (Kim et al., 2023) distilled social interactions dialogue by contextualizing commonsense knowledge from a social knowledge graph. PLACES (Chen et al., 2023a) generated a multi-party social dialog by prompting the OPT (Zhang et al., 2022) model with topic, facts, and dialog as in-context learning examples. MathDial (Macina et al., 2023) distilled teacher-student math reasoning tutoring dialog through pairing human teachers with a LLM. There are also works that distil multi-turn open-domain instructional conversations, such as Baize (Xu et al., 2023) and UltraChat (Ding et al., 2023). Our TOAD represents the first work in distilling TOD dialog with multiple response style options, which requires more nuanced control over the conversation content.
@@ 3.1 Persona-Grounded Context Generation
Data diversity is important to the model’s robustness and generalization. However, LLM-generated data given the same input prompt often lacks diversity. To address the two issues mentioned above, we develop a persona-grounded context generation pipeline, which combines sampling from external data sources and chain-of-thought generation (Wei et al., 2022). Recent work (Hu and Collier, 2024) has shown that LLMs are capable to simulate different perspectives.
@@
  Compound restaurant_booking_reserve_table(restaurant ="French Brasserie", time="8:00 PM"), I would like to reserve a table at the French Brasserie for 8:00 PM. Also, can you help me find a hotel in Las Vegas?
@@
  Compositional weather_get_weather(date=get_calendar_events( name="Art Class").calendar_events_check(date).date) What’s the weather like on the day of my Art Class event?
  Self-correction get_movie_time(movie_name="Fast & Furious Presents: Hobbs & Shaw", location="Miami").self_correction( location="Houston") Could you find the showtimes for Fast & Furious Presents: Hobbs & Shaw in Miami? Actually, make that Houston instead.
  Complex referral get_alarms(ordered_by="time", index=0).check(time) What’s the time for my earliest alarm?
@@ 3.4 Dataset Quality Control
To ensure the quality of the generated dataset, we conducted a two-step quality control process. The first step involved two automatic sanity checks: 1) Removing datapoints with utterances containing unformatted values, such as internal datetime representations like ‘2025-12-16’. 2) Removing datapoints that have missing slot values in the corresponding utterances.
The second step addressed the potential inconsistencies between plot actions and utterances that may arise due to the non-deterministic nature of language models. Although human evaluation is considered the gold standard, it is not always feasible or scalable for large-scale dataset expansion. As an alternative, we propose using a language model evaluator to perform quality control, an approach that has gained popularity in recent works (Chen et al., 2023b; Zeng et al., 2023; Liu et al., 2024). These studies have demonstrated the effectiveness of language model evaluators in assessing various aspects of generation tasks, such as summarization and data-to-text.
To implement this approach, we provide the LLM evaluator (gpt-3.5-turbo) with evaluation criteria, app context, action plot, and corresponding utterances. The evaluator then determines whether the utterances are consistent with the action plot and the app context. In our case, the LLM labeled 37 data points as inconsistent. Upon manual examination, we confirmed that 32 of these samples were indeed inconsistent. We listed the template used for quality control in Appendix I.
@@ 9 Ethics and Risks
To ensure no user is disadvantaged, a dataset for model training needs to represent a diverse range of users. The distribution of persona attributes in the TOAD dataset is detailed in Appendix E.
@@ Appendix E Persona and App Context Statistics
In TOAD, we simulate 500 personas with diverse backgrounds, with one persona randomly assigned to each dialog as the user’s information source. The gender distribution of the simulated personas is approximately 52% male and 48% female.
@@ Appendix G Plot and Dialog Examples
{’today’:’Wed 2025-04-16’, ’alarms’:[{’time’:’07:00’, ’name’:’Morning Workout’, ’if_repeat’:True}, {’time’:’18:30’, ’name’:’Family Dinner’, ’if_repeat’:True}]} User@: [get_alarms(ordered_by="time", index=0).alarms_delete(name), get_alarms(ordered_by="time", index=1).alarms_check(name)] System@: [alarms_ask_confirm(get_alarms(name="Morning Workout").delete(name)), alarms_inform_result(name="Family Dinner")] User@: alarms_confirm() System@: alarms_notify_done(get_alarms(name="Morning Workout").delete(name)) User@: Can you delete my earliest alarm and check the next one? System@: { HV M: Are you sure you want to delete the Morning Workout alarm? Your next alarm is Family Dinner at 18:30. HV no_M: Do you confirm the deletion of your earliest alarm, Morning Workout’? The following alarm is Family Dinner at 18:30. MV M: You want to delete the Morning Workout alarm, correct? Your next one is Family Dinner at 18:30. MV no_M: Confirm to delete this alarm? The next one is at 18:30 for Family Dinner’. LV M: Delete Morning Workout’? Next: Family Dinner 18:30. LV no_M: Confirm deletion? Next: 18:30 Family Dinner’. } User@: Yes, go ahead and delete it. System@: { HV M: Your Morning Workout alarm has been deleted. HV no_M: The earliest alarm has been deleted. MV M: Morning Workout alarm is deleted. MV no_M: Its deleted. LV M: Deleted. LV no_M: Done.}
{’today’:’Mon 2026-12-28’} User@: purchase_tickets(movie_name="Joker", date="Tue 2026-12-29", cinema_name="Regal Cinemas", ticket_quantity=1, movie_format="3d") System@: request_infomation(showtime) User@: inform_infomation(showtime="20:00") System@: ask_confirm(purchase_tickets(movie_name="Joker", ticket_quantity=1, date="Tue 2026-12-29", showtime="20:00", movie_format="3d", cinema_name="Regal Cinemas")) User@: confirm() System@: notify_done(purchase_tickets(movie_name="Joker", ticket_quantity=1, date="Tue 2026-12-29", showtime="20:00", movie_format="3d", cinema_name="Regal Cinemas")) User@: I’d like to buy a 3D ticket for Joker at Regal Cinemas tomorrow, please. System@: { HV M: What time would you like to watch Joker at Regal Cinemas tomorrow? HV no_M: What showtime do you prefer? MV M: When would you like to see it? MV no_M: When is the preferred time? LV M: Showtime? LV no_M: Time? } User@: At 20:00. System@: { HV M: So, youd like to purchase a 3D ticket for Joker at Regal Cinemas on Tue 2026-12-29 at 20:00. Is that correct? HV no_M: Just to confirm, you want to buy a 3D ticket for the movie Joker at Regal Cinemas on Tue 2026-12-29 at 20:00, right? MV M: Youre buying a 3D ticket for Joker at Regal Cinemas tomorrow at 20:00. Correct? MV no_M: So, youre purchasing a ticket for Joker in 3D at Regal Cinemas tomorrow at 20:00. Is that right? LV M: Confirm purchase? LV no_M: Confirm? } User@: Yes, that’s correct. System@: { HV M: Your 3D ticket for Joker at Regal Cinemas on Tue 2026-12-29 at 20:00 has been purchased. HV no_M: The ticket has been successfully purchased. MV M: Ive bought your ticket for that movie at that time. MV no_M: Its done, the ticket is bought. LV M: Done. LV no_M: Purchased. }
@@ Appendix H Dialog Generator Template
Prompt template for user utterance generation
"""\ You are a smartphone user and you are testing your virtual assistant on your phone by \ engaging in a multi-turn conversations with it. Here is your personal introduction: {{ user_intro }} Instructions: 1. You need to communicate with the assistant following the guidance of "actions". 2. Based on the introduction, think about what speech habit you should have and \ communicate with this pattern. 3. The "message" should have minimum words possible. 4. You must return in JSON format, following the provided examples. The context information is: {{ context }}. You are using these apps: {{ situation }}. Begin conversation (you are identified as "user"). user: {"actions": ["hello()"], "message": "Hi."} assistant: {"actions": ["offer_help()"], "message": "Hello, how can I help?"} {if dialog_history | length>0 %}{{ dialog_history }}{endif %} The "actions" you need to follow is {{ action[cur_turn] }}. What are you going to say next? \ {{ user_style_instruction }} user:\ """
@@
"""\ Instructions: 1. You are a virtual assistant. Your goal is to assist the user to accomplish their goal. 2. Your responses should strictly follow the given actions and be helpful, natural, \ professional and concise. 3. Your response should strictly follow the corresponding "actions". The user is interacting with these apps: {{ situation }}. Style instruction: verbosity_low’: Your response must only have a couple of words, such as "when", \ "how long" or "done". verbosity_mid’: Your response should be a concise but complete sentence and must \ replace the nouns or noun phrases mentioned by user with pronouns, such as "it", "that" and "its". verbosity_high’: Your response should use full expressions with all the details. mirroring’: Your response should use the users noun phrase or verb expressions when possible. no_mirroring’: Ignore all previous dialog. Do not affect by user expression. summary’: Your response should be a brief report of the given summary. Response Example: assistant actions: {"verbosity_low mirroring": ["notify_done()"], \ "verbosity_low no_mirroring": ["notify_done()"], \ "verbosity_mid mirroring": ["notify_done(operation_on_device(\ operation="turn_off", device="heating", home_space="bedroom"))"], \ "verbosity_mid no_mirroring": ["notify_done(operation_on_device(\ operation="turn_off", device="heating", home_space="bedroom"))"], \ "verbosity_high mirroring": ["notify_done(operation_on_device(\ operation="turn_off", device="heating", home_space="bedroom"))"], \ "verbosity_high no_mirroring": ["notify_done(operation_on_device(\ operation="turn_off", device="heating", home_space="bedroom"))"]} assistant: {"verbosity_low mirroring": "Turned off.", \ "verbosity_low no_mirroring": "Done.", \ "verbosity_mid mirroring": "I have turned off the heating in that room.", \ "verbosity_mid no_mirroring": "I have turned it off in that room.", \ "verbosity_high mirroring": "I have turned off the bedroom heating.", \ "verbosity_high no_mirroring": "Sure, I have turned off the heating in the bedroom."} Conversation history: {if dialog_history | length>0 %}{{ dialog_history }}{endif %} New turn: You should return in JSON format with 6 keys: ["verbosity_low mirroring", \ "verbosity_low no_mirroring", "verbosity_mid mirroring", "verbosity_mid no_mirroring", \ "verbosity_high mirroring", "verbosity_high no_mirroring"]. assistant actions: {{ action[cur_turn] }}. assistant: \ """
@@
Appendix I Quality Control Template
Prompt template for quality control
"""\ Please check the consistency between the actions and the corresponding utterances for the following dialog. They might refer to the context below. Context: {{ context }} Dialog: {{ actions_and_utterances }} Response: \ """
Table 8: Prompt template example for examining the consistency between action plot and utterances